# Watson-Crick Conjugates of Words and Languages

### v3

@@ Abstract.
Combinatorics on words; conjugate; Watson-Crick conjugate; antimorphic involution; Theoretical DNA computing
1. Introduction
@@
Conjugates of a word are significant in the field of combinatorics on words as they have numerous applications in cryptography ([3]), biology ([38]), and other areas. Our analysis in this work focuses on various combinatorial, structural, and language-theoretic properties of Watson-Crick conjugates, or $\theta$-conjugates (where $\theta$ is an antimorphic involution) of words - a generalization of the conjugates of words.
@@
Over the years, DNA sequences have been used to encode/design a range of complex problems/experiments (for example, see [1], [2], [4], [5], [11], [12], [13], [14], [16], [31], [34], [36], [39], [40], etc). When encoding or designing a problem using DNA sequences, the occurrence of inter/intra-molecular hybridizations (see Fig. 1) between DNA strands can be perceived both as an advantage and a disadvantage. For example, formation of hairpin structures have found utility in various applications, including solving the satisfiability problem ([34]), designing DNA aptamers ([14]), and programming catalytic DNA self-assembly strategies ([4]). Conversely, these hybridizations are sometimes viewed as a disadvantage (see [36, 40]). This is primarily because these hybridizations can significantly disrupt the intended interactions between the involved sequences in inter/intra-molecular hybridizations and other DNA strands in the preprogrammed and expected ways. Therefore, to avoid such undesirable hybridizations, careful DNA sequence design is necessary at the time of encoding the problem. Drawing inspiration from this approach of DNA sequence design, several authors have proposed various conditions for encoding by extending various concepts of classical combinatorics on words, such as primitive words ([7], [8], [10], [17], [24]), bordered words ([18], [19], [20], [21], [22], [25]), palindromes ([22], [23]), and rich words ([37]).
$y$$\theta_{WC}(y)$$\beta$$\gamma$$x$$\theta_{WC}(x)$$\alpha$
Figure 1. Example of inter- and intra- molecular structures
@@
In [20], the authors extended the concept of bordered word to $\theta$-bordered word from the perspective of DNA computing, where $\theta$ is an antimorphic involution (i.e., $\theta(xy)=\theta(y)\theta(x)$ and $\theta^{2}(t)=t$ for any $x,y,t\in\Sigma^{*}$) that incorporates the notion of the Watson-Crick complementarity of DNA molecules. They also defined $\theta$-bordered word for a morphic involution $\theta$ (i.e., $\theta(xy)=\theta(x)\theta(y)$ and $\theta^{2}(t)=t$ for any $x,y,t\in\Sigma^{*}$) as a generalization of bordered word. For a morphic/antimorphic involution $\theta$, a word $w$ is said to be $\theta$-bordered if $w=uz=\theta(z)v$ for some non-empty words $u,z,v$. A word that is not $\theta$-bordered is called $\theta$-unbordered. When focusing on the DNA alphabet $\{A,T,G,C\}$ and utilizing the Watson-Crick complementarity function $\theta_{WC}$ ($\theta_{WC}$ is an antimorphic involution on $\{A,T,G,C\}^{*}$ such that $\theta_{WC}(A)=T$, $\theta_{WC}(T)=A$, $\theta_{WC}(C)=G$, $\theta_{WC}(G)=C$), $\theta_{WC}$-unbordered words represents the collection of DNA single strands that are free from both inter- and intra-molecular structures such as the ones shown in Fig. 1(left), and hairpins shown in Fig. 1(right). From the definition of $\theta$-bordered word, authors of [22] defined the notion of $\theta$-conjugate of a word for morphic/antimorphic involution $\theta$. Formally, a word $u$ is called a $\theta$-conjugate of a word $v$ if there exists a word $z$ such that $uz=\theta(z)v$. The set of all $\theta$-conjugates of $v$ is denoted by $C_{\theta}(v)$. When $\theta$ is an antimorphic involution over $\Sigma^{*}$, then analyzing the solution of the equation $uz=\theta(z)v$ (as presented in Proposition $2$ in [22]), it becomes eviden that for $v=xy$, the $\theta$-conjugate $u$ can be expressed as $\theta(y)x$ for some $x,y\in\Sigma^{*}$ (see Fig. 2).
Figure 2. Structure of conjugates &\&& $\theta$-conjugates of word for antimorphic involution $\theta$
$x$$y$$\theta_{WC}(y)$$x$
@@
Figure 3. Inter-molecular structure between a word and its $\theta_{WC}$-conjugate
@@
In [29], authors discussed several properties of the set of all $\theta$-conjugates of a word when $\theta$ is an antimorphic involution. They mainly studied the cardinality of the set of all $\theta$-conjugates of a word and occurrences of palindromes, $\theta$-palindromes in the set of all $\theta$-conjugates of a word. In this work, we further explore properties of the set of all $\theta$-conjugates of a word when $\theta$ is an antimorphic involution. We also extend the idea of $\theta$-conjugates of a word to $\theta$-conjugates of a language $L$ as $C_{\theta}(L)=\bigcup\limits_{w\in L}C_{\theta}(w)$ and study various properties of $C_{\theta}(L)$ when $\theta$ is an antimorphic involution. The text is organized as follows. In Section 3, we study some properties of the set of all $\theta$-conjugates of a word. We first investigate the structure of $\theta$-conjugates of a word. Then, we discuss the solution of the equation $C_{\theta}(u)=C_{\theta}(v)$. Section 4 investigates the Chomsky hierarchy for $C_{\theta}(L)$ when $L$ is regular, context-free, context-sensitive and recursively enumerable. In Section 5, we define and discuss the iterated $\theta$-conjugate of words and languages. Section 6 investigates some decidability problems for $\theta$-conjugate-free languages. We end our work with some concluding remarks.
2. Basic definitions and notations
@@
An alphabet $\Sigma$ is a finite non-empty set of letters. A word over $\Sigma$ is a sequence of letters from $\Sigma$. The length of a word $w$, denoted by $|w|$, is the number of letters in $w$. A word with length zero is called the empty word and is denoted by $\lambda$. The set of all finite words over $\Sigma$ is denoted by $\Sigma^{*}$, and $\Sigma^{+}=\Sigma^{*}\setminus\{\lambda\}$. A language $L$ over $\Sigma$ is a subset of $\Sigma^{*}$. The set of all words over $\Sigma$ of length $n$ is denoted by $\Sigma^{n}$ and the set of all words over $\Sigma$ of length greater than or equal to $n$ is denoted by $\Sigma^{\geq n}$. Let $\operatorname{Alph}(w)$ denote the set of all letters that occur in a word $w$ and $\operatorname{Alph}(L)=\bigcup\limits_{w\in L}\operatorname{Alph}(w)$ where $L$ is a language. A word $u$ is a factor of a word $w$ if $w=xuy$ for some $x,y\in\Sigma^{*}$. When $x=\lambda$ (respectively $y=\lambda$), $u$ is called a prefix (respectively suffix) of $w$. If $u$ is a prefix (respectively suffix) of $w$ and $u\neq w$, then $u$ is called a proper prefix (respectively proper suffix) of $w$. Set of all factors (respectively prefixes and suffixes) of a word $w$ is denoted by $Fac(w)$ (respectively $Pref(w)$ and $Suf(w)$). For $x\in\Sigma^{*}$, $|w|_{x}$ denotes the number of occurrences of $x$ in $w$. For a word $w=w_{1}w_{2}\cdots w_{n}$ where each $w_{i}\in\Sigma$, the reversal of $w$ is $w^{R}=w_{n}\cdots w_{2}w_{1}$. A word $w$ is a palindrome if $w=w^{R}$. A word $u$ is a conjugate of a word $w$ if $u=yx$ and $w=xy$ for some $x,y\in\Sigma^{*}$ (see Figure 2). The set of all conjugates of $w$ is denoted by $C(w)$. A word $w\in\Sigma^{+}$ is called primitive if $w=x^{n}$ implies $w=x$ and $n=1$. For every $w\in\Sigma^{+}$, there exists a unique $\rho(w)\in\Sigma^{+}$, called the root of $w$, such that $\rho(w)$ is a primitive word and $w=\rho(w)^{n}$ for some integer $n\geq 1$.
@@
A deterministic finite automaton (DFA) $M$ is defined by $M=(S^{\prime},\Sigma,\delta,q_{0},F)$, where $S^{\prime}$ is a finite set of states, $\Sigma$ is an alphabet, $\delta:S^{\prime}\times\Sigma\rightarrow S^{\prime}$ is a transition function, $q_{0}\in S^{\prime}$ is the initial state and $F\subseteq S^{\prime}$ is the set of all final states. A word is accepted by the DFA $M$ if, starting from the initial state $q_{0}$, processing the word using the transition function $\delta$ leads to a state in $F$. The set of all words accepted by $M$ is denoted by $L(M)$. A language $L_{1}$ is said to be regular if there exists a DFA $M_{1}$ such that $L_{1}=L(M_{1})$. A linear bounded automaton is a non-deterministic Turing machine $M^{\prime}=(Q^{\prime},\Sigma,\Gamma,\delta,q_{0},<,>,B,F)$ where $Q^{\prime}$ is the set of all states, $\Gamma$ is the tape alphabet, $\Sigma\subseteq\Gamma\setminus\{B\}$ is the input alphabet, $\delta:Q^{\prime}\times\Gamma\rightarrow 2^{Q^{\prime}\times\Gamma\times\{L,R,N\}}$ is the transition function, $B\in\Gamma$ is the blank symbol, $q_{0}\in Q^{\prime}$ is the initial state, $F\subseteq Q^{\prime}$ is the set of final states, $<\in\Sigma$ is the left end marker and $>\in\Sigma$ is the right end marker. Here, $L,R$ and $N$ denote the possible movements of the tape head: left, right, and no movement, respectively, as dictated by the transition function $\delta$. A word $w$ is accepted by the linear bounded automaton $M^{\prime}$ if the automaton, starting in its initial state $q_{0}$ with the input word $<w>$ on the tape, can transition through a sequence of states according to its transition function $\delta$ and eventually reach one of its final states in $F$. Here note that if $w$ is a word accepted by a linear bounded automaton $M^{\prime}$, then $w$ is in $(\Sigma\setminus\{<,>\})^{*}$. For a state $q$ and two strings $u$, $v$ over the tape alphabet $\Gamma$, we write $uqv$ for the configuration of the Turing machine $M^{\prime}$ where the current state is $q$, the current tape content is $uv$, and the current head location is the first symbol of $v$. A move of the Turing machine $M^{\prime}$ from one configuration to another is denoted by $\vdash_{M^{\prime}}$. The symbol ${\vdash}^{*}_{M^{\prime}}$ represents an arbitrary number of moves of $M^{\prime}$.
For all other concepts in formal language theory and combinatorics on words, the reader is referred to [26, 27, 33, 35].
@@
[22] For $u,v\in\Sigma^{+}$, if $uv=\theta(v)u$ and $\theta$ is an antimorphic involution, then $u=x(yx)^{i}$, $v=yx$ for some integer $i\geq 0$ and $\theta$-palindromes $x\in\Sigma^{+}$, $y\in\Sigma^{*}$.
3. Some Important Properties of set of all $\theta$-conjugates of a Word
@@
To discuss the structure of elements of $C_{\theta}(w)$ for some $w$, we first recall the following result from [22].
@@
[22] For an antimorphic involution $\theta$, consider $u$ is a $\theta$-conjugate of $w$, i.e., $uv=\theta(v)w$ for some $v\in\Sigma^{*}$. Then, either $u=xy$, $v=\theta(x)$ and $w=y\theta(x)$ for some $x,y\in\Sigma^{*}$, or $w=\theta(u)$ and $v=\beta w$ for some $\theta$-palindrome $\beta$.
@@
In the following, we illustrate Proposition 3.2 with several examples. These examples also highlight an important property of the $\theta$-conjugacy relation. Similar to the conjugacy relation, we define the $\theta$-conjugacy relation $R_{\theta}$ on $\Sigma^{*}$, where $(u,v)\in R_{\theta}$ indicates that $u$ is a $\theta$-conjugate of $v$. It is known that the conjugacy relation is an equivalence relation [27]. However, with the help of the following examples, we show that the $\theta$-conjugacy relation is not an equivalence relation.
@@
By definition of conjugates, if $w\in\Sigma^{+}$ and $a\in\operatorname{Alph}(w)$, then there exists an element in $C(w)$ that begins with $a$. However, this property does not hold for $\theta$-conjugates, as illustrated in Example 3.3. In the following, we discuss the structure of elements of the set of all $\theta$-conjugates of a word regarding this property.
@@
  1. (1)
    If $w=bccb$, then $C_{\theta}(w)=\{bccb,abcc,acbc,accb,acca\}$. Note that $bccb$ is not a $\theta$-palindrome. Now, $C_{\theta}(bcc)=\{bcc,cbc,ccb,cca\}$ and $C_{\theta}(bccb)=\{bccb\}\cup\theta(b)C_{\theta}(bcc)$.
  2. (2)
@@
For $w\in\Sigma^{+}$, $C(w)$ can contain at most $|w|$ elements and if $|C(w)|=|w|$, then each element of $C(w)$ is primitive. Authors in [29] proved that the set of all $\theta$-conjugates of a word $w$ can have at most $|w|+1$ elements. We now prove that for $w\in\Sigma^{*}$, if $|C_{\theta}(w)|=|w|+1$, then $C_{\theta}(w)$ must contain at least one primitive word. To prove this we need the following result.
@@
  • For $n\geq 2$, let us assume that each element of $C_{\theta}(w)$ is non-primitive. Now, the length of the root of each element in $C_{\theta}(w)$ is at most $\lfloor{\frac{|w|}{2}}\rfloor$. Consider $l=\lfloor{\frac{|w|}{2}}\rfloor$ and $w=w_{1}w_{2}\cdots w_{l}w_{l+1}\cdots w_{n}$ where each $w_{j}\in\Sigma$ for $1\leq j\leq n$. For each $i~{}(0\leq i\leq l-1)$, let $\alpha_{i}=\beta_{i}\gamma_{i}$ where $\beta_{i}=\theta(w_{l-i}\cdots w_{n})$ and $\gamma_{i}=w_{1}w_{2}\cdots w_{l-i-1}$. Then for each $i~{}(0\leq i\leq l-1)$, $|\beta_{i}|>l$ and $\alpha_{i}\in C_{\theta}(w)$. Consider $\rho(\alpha_{i})=z_{i}$ for each $i~{}(0\leq i\leq l-1)$. Then, $|z_{i}|\leq l$ for each $i~{}(0\leq i\leq l-1)$. Now, each $\alpha_{i}~{}(0\leq i\leq l-1)$ has a common prefix $\beta_{0}=\theta(w_{n})\theta(w_{n-1})\cdots\theta(w_{l})$ of length greater than $l$. This implies each $z_{i}$ is a proper prefix of $\beta_{0}$ as $|z_{i}|\leq l$ where $10\leq i\leq l-$ 1. Since each element of $C_{\theta}(w)$ is non-primitive and $|C_{\theta}(w)|=|w|+1$, each $z_{i}$  $(0\leq i\leq l-1)$ is a distinct prefix of $\beta_{0}$. Then, the possible lengths of $z_{i}$’s are $1,2,\cdots,l$.
@@
    Otherwise, consider the length of each $z_{i}$ is distinct. Since the length of the root of each element in $C_{\theta}(w)$ is at most $l$ and there are $l$ distinct elements $\alpha_{0},\alpha_{1},\ldots,\alpha_{l-1}$ in $C_{\theta}(w)$ with a commom prefix $\beta_{0}$, we have $|\rho(\alpha_{i_{3}})|=1$ for some $10\leq i_{3}\leq l-$ 1. Then by Lemma 3.8, $w=a^{n}$ with $\theta(a)=a$, which is a contradiction to our assumption that $|C_{\theta}(w)|=|w|+1$.
@@
For $u,v\in\Sigma^{*}$, the equality of $C_{\theta}(u)$ and $C_{\theta}(v)$ was studied in [29], and the following result was proved.
@@
[29] Let $u,v\in\Sigma^{*}$ such that $C_{\theta}(u)=C_{\theta}(v)$.
  1. (1)
@@
    2. (ii)
      $u=a^{n_{1}}u^{\prime}a^{n_{2}}$ and $v=u^{R}$ where $a\in\Sigma$, $u^{\prime}\in\Sigma^{+}$, $a\notin Pref(u^{\prime})$, $a\notin Suf(u^{\prime})$, $u^{\prime}=u^{\prime R}$, $n_{1},n_{2}\geq 1$, $n_{1}\neq n_{2}$, and $\theta(c)=c~{}\forall c\in\operatorname{Alph}(u)$.

### v5

@@ Abstract
Combinatorics on words, conjugate, Watson-Crick conjugate, antimorphic involution, Theoretical DNA computing
1 Introduction
@@
Conjugates of a word are significant in the field of combinatorics on words as they have numerous applications in cryptography (Belkheyar et al. (2023)), biology (Walter et al. (2015)), and other areas. Our analysis in this work focuses on various combinatorial, structural, and language-theoretic properties of Watson-Crick conjugates, or $\theta$-conjugates (where $\theta$ is an antimorphic involution) of words - a generalization of the conjugates of words.
@@
Over the years, DNA sequences have been used to encode/design a range of complex problems/experim-ents (for example, see Adleman (1994), Aoi et al. (1998), Bi et al. (2016), Braich et al. (2002), Head et al. (2000), Henkel et al. (2007), Henkel et al. (2005), Jeddi and Saiz (2017), Johnson (2008), Ouyang et al. (1997), Sakamoto et al. (2000), Shi et al. (2017), Wang et al. (2008), Zhang and Seelig (2011), etc). When encoding or designing a problem using DNA sequences, the occurrence of inter/intra-molecular hybridizations (see Fig. 1) between DNA strands can be perceived both as an advantage and a disadvantage. For example, formation of hairpin structures have found utility in various applications, including solving the satisfiability problem (Sakamoto et al. (2000)), designing DNA aptamers (Jeddi and Saiz (2017)), and programming catalytic DNA self-assembly strategies (Bi et al. (2016)). Conversely, these hybridizations are sometimes viewed as a disadvantage (see Shi et al. (2017); Zhang and Seelig (2011)). This is primarily because these hybridizations can significantly disrupt the intended interactions between the involved sequences in inter/intra-molecular hybridizations and other DNA strands in the preprogrammed and expected ways. Therefore, to avoid such undesirable hybridizations, careful DNA sequence design is necessary at the time of encoding the problem. Drawing inspiration from this approach of DNA sequence design, several authors have proposed various conditions for encoding by extending various concepts of classical combinatorics on words, such as primitive words (Chiniforooshan et al. (2010), Czeizler et al. (2010), Gawrychowski et al. (2014), Kari and Kulkarni (2014a), Kari and Ng (2019)), bordered words (Kari and Kulkarni (2014b), Kari and Kulkarni (2017), Kari and Mahalingam (2007), Kari and Mahalingam (2008a), Kari and Mahalingam (2008b), Kari and Seki (2009)), palindromes (Kari and Mahalingam (2008b), Kari and Mahalingam (2010)), and rich words (Starosta (2011)).
[…]
Figure 1: Example of inter- and intra- molecular structures
@@
In Kari and Mahalingam (2007), the authors extended the concept of bordered word to $\theta$-bordered word from the perspective of DNA computing, where $\theta$ is an antimorphic involution (i.e., $\theta(xy)=\theta(y)\theta(x)$ and $\theta^{2}(t)=t$ for any $x,y,t\in\Sigma^{*}$) that incorporates the notion of the Watson-Crick complementarity of DNA molecules. They also defined $\theta$-bordered word for a morphic involution $\theta$ (i.e., $\theta(xy)=\theta(x)\theta(y)$ and $\theta^{2}(t)=t$ for any $x,y,t\in\Sigma^{*}$) as a generalization of bordered word. For a morphic/antimorphic involution $\theta$, a word $w$ is said to be $\theta$-bordered if $w=uz=\theta(z)v$ for some non-empty words $u,z,v$. A word that is not $\theta$-bordered is called $\theta$-unbordered. When focusing on the DNA alphabet $\{A,T,G,C\}$ and utilizing the Watson-Crick complementarity function $\theta_{WC}$ ($\theta_{WC}$ is an antimorphic involution on $\{A,T,G,C\}^{*}$ such that $\theta_{WC}(A)=T$, $\theta_{WC}(T)=A$, $\theta_{WC}(C)=G$, $\theta_{WC}(G)=C$), $\theta_{WC}$-unbordered words represents the collection of DNA single strands that are free from both inter- and intra-molecular structures such as the ones shown in Fig. 1(left), and hairpins shown in Fig. 1(right). From the definition of $\theta$-bordered word, Kari and Mahalingam (2008b) defined the notion of $\theta$-conjugate of a word for morphic/antimorphic involution $\theta$. Formally, a word $u$ is called a $\theta$-conjugate of a word $v$ if there exists a word $z$ such that $uz=\theta(z)v$. The set of all $\theta$-conjugates of $v$ is denoted by $C_{\theta}(v)$. When $\theta$ is an antimorphic involution over $\Sigma^{*}$, then analyzing the solution of the equation $uz=\theta(z)v$ (as presented in Proposition $2$ in Kari and Mahalingam (2008b)), it becomes evident that for $v=xy$, the $\theta$-conjugate $u$ can be expressed as $\theta(y)x$ for some $x,y\in\Sigma^{*}$ (see Fig. 2).
Figure 2: Structure of conjugates &\& $\theta$-conjugates of word for antimorphic involution $\theta$
[…]
@@
Figure 3: Inter-molecular structure between a word and its $\theta_{WC}$-conjugate
@@
Mahalingam et al. (2022) discussed several properties of the set of all $\theta$-conjugates of a word when $\theta$ is an antimorphic involution. They mainly studied the cardinality of the set of all $\theta$-conjugates of a word and occurrences of palindromes, $\theta$-palindromes in the set of all $\theta$-conjugates of a word. In this work, we further explore properties of the set of all $\theta$-conjugates of a word when $\theta$ is an antimorphic involution. We also extend the idea of $\theta$-conjugates of a word to $\theta$-conjugates of a language $L$ as $C_{\theta}(L)=\bigcup\limits_{w\in L}C_{\theta}(w)$ and study various properties of $C_{\theta}(L)$ when $\theta$ is an antimorphic involution. The text is organized as follows. In Section 3, we study some properties of the set of all $\theta$-conjugates of a word. We first investigate the structure of $\theta$-conjugates of a word. Then, we discuss the solution of the equation $C_{\theta}(u)=C_{\theta}(v)$. Section 4 investigates the Chomsky hierarchy for $C_{\theta}(L)$ when $L$ is regular, context-free, context-sensitive and recursively enumerable. In Section 5, we define and discuss the iterated $\theta$-conjugate of words and languages. Section 6 investigates some decidability problems for $\theta$-conjugate-free languages. We end our work with some concluding remarks.
2 Basic definitions and notations
@@
An alphabet $\Sigma$ is a finite non-empty set of letters. A word over $\Sigma$ is a sequence of letters from $\Sigma$. The length of a word $w$, denoted by $|w|$, is the number of letters in $w$. A word with length zero is called the empty word and is denoted by $\lambda$. The set of all finite words over $\Sigma$ is denoted by $\Sigma^{*}$, and $\Sigma^{+}=\Sigma^{*}\setminus\{\lambda\}$. A language $L$ over $\Sigma$ is a subset of $\Sigma^{*}$. The set of all words over $\Sigma$ of length $n$ is denoted by $\Sigma^{n}$ and the set of all words over $\Sigma$ of length greater than or equal to $n$ is denoted by $\Sigma^{\geq n}$. Let $\textit{Alph}(w)$ denote the set of all letters that occur in a word $w$ and $\textit{Alph}(L)=\bigcup\limits_{w\in L}\textit{Alph}(w)$ where $L$ is a language. A word $u$ is a factor of a word $w$ if $w=xuy$ for some $x,y\in\Sigma^{*}$. When $x=\lambda$ (respectively $y=\lambda$), $u$ is called a prefix (respectively suffix) of $w$. If $u$ is a prefix (respectively suffix) of $w$ and $u\neq w$, then $u$ is called a proper prefix (respectively proper suffix) of $w$. Set of all factors (respectively prefixes and suffixes) of a word $w$ is denoted by $Fac(w)$ (respectively $Pref(w)$ and $Suf(w)$). For $x\in\Sigma^{*}$, $|w|_{x}$ denotes the number of occurrences of $x$ in $w$. For a word $w=w_{1}w_{2}\cdots w_{n}$ where each $w_{i}\in\Sigma$, the reversal of $w$ is $w^{R}=w_{n}\cdots w_{2}w_{1}$. A word $w$ is a palindrome if $w=w^{R}$. A word $u$ is a conjugate of a word $w$ if $u=yx$ and $w=xy$ for some $x,y\in\Sigma^{*}$ (see Figure 2). The set of all conjugates of $w$ is denoted by $C(w)$. A word $w\in\Sigma^{+}$ is called primitive if $w=x^{n}$ implies $w=x$ and $n=1$. For every $w\in\Sigma^{+}$, there exists a unique $\rho(w)\in\Sigma^{+}$, called the root of $w$, such that $\rho(w)$ is a primitive word and $w=\rho(w)^{n}$ for some integer $n\geq 1$.
@@
A deterministic finite automaton (DFA) $M$ is defined by $M=(S^{\prime},\Sigma,\delta,q_{0},F)$, where $S^{\prime}$ is a finite set of states, $\Sigma$ is an alphabet, $\delta:S^{\prime}\times\Sigma\rightarrow S^{\prime}$ is a transition function, $q_{0}\in S^{\prime}$ is the initial state and $F\subseteq S^{\prime}$ is the set of all final states. A word is accepted by the DFA $M$ if, starting from the initial state $q_{0}$, processing the word using the transition function $\delta$ leads to a state in $F$. The set of all words accepted by $M$ is denoted by $L(M)$. A language $L_{1}$ is said to be regular if there exists a DFA $M_{1}$ such that $L_{1}=L(M_{1})$. A linear bounded automaton is a non-deterministic Turing machine $M^{\prime}=(Q^{\prime},\Sigma,\Gamma,\delta,q_{0},<,>,B,F)$ where $Q^{\prime}$ is the set of all states, $\Gamma$ is the tape alphabet, $\Sigma\subseteq\Gamma\setminus\{B\}$ is the input alphabet, $\delta:Q^{\prime}\times\Gamma\rightarrow 2^{Q^{\prime}\times\Gamma\times\{L,R,N\}}$ is the transition function, $B\in\Gamma$ is the blank symbol, $q_{0}\in Q^{\prime}$ is the initial state, $F\subseteq Q^{\prime}$ is the set of final states, $\in\Sigma$ is the left end marker and $\in\Sigma$ is the right end marker. Here, $L,R$ and $N$ denote the possible movements of the tape head: left, right, and no movement, respectively, as dictated by the transition function $\delta$. A word $w$ is accepted by the linear bounded automaton $M^{\prime}$ if the automaton, starting in its initial state $q_{0}$ with the input word <w><w> on the tape, can transition through a sequence of states according to its transition function $\delta$ and eventually reach one of its final states in $F$. Here note that if $w$ is a word accepted by a linear bounded automaton $M^{\prime}$, then $w$ is in $(\Sigma\setminus\{<,>\})^{*}$. For a state $q$ and two strings $u$, $v$ over the tape alphabet $\Gamma$, we write uqvuqv for the configuration of the Turing machine $M^{\prime}$ where the current state is $q$, the current tape content is $uv$, and the current head location is the first symbol of $v$. A move of the Turing machine $M^{\prime}$ from one configuration to another is denoted by $\vdash_{M^{\prime}}$. The symbol ${\vdash}^{*}_{M^{\prime}}$ represents an arbitrary number of moves of $M^{\prime}$.
For all other concepts in formal language theory and combinatorics on words, the reader is referred to Linz (2012); Lothaire (1997); Rozenberg and Salomaa (Eds.); Shallit (2008).
@@
Kari and Mahalingam (2008b) For $u,v\in\Sigma^{+}$, if $uv=\theta(v)u$ and $\theta$ is an antimorphic involution, then $u=x(yx)^{i}$, $v=yx$ for some integer $i\geq 0$ and $\theta$-palindromes $x\in\Sigma^{+}$, $y\in\Sigma^{*}$.
3 Some Important Properties of set of all $\theta$-conjugates of a Word
@@
To discuss the structure of elements of $C_{\theta}(w)$ for some $w$, we first recall the following result from Kari and Mahalingam (2008b).
@@
Kari and Mahalingam (2008b) For an antimorphic involution $\theta$, consider $u$ is a $\theta$-conjugate of $w$, i.e., $uv=\theta(v)w$ for some $v\in\Sigma^{*}$. Then, either $u=xy$, $v=\theta(x)$ and $w=y\theta(x)$ for some $x,y\in\Sigma^{*}$, or $w=\theta(u)$ and $v=\beta w$ for some $\theta$-palindrome $\beta$.
@@
In the following, we illustrate Proposition 3.2 with several examples. These examples also highlight an important property of the $\theta$-conjugacy relation. Similar to the conjugacy relation, we define the $\theta$-conjugacy relation $R_{\theta}$ on $\Sigma^{*}$, where $(u,v)\in R_{\theta}$ indicates that $u$ is a $\theta$-conjugate of $v$. It is known that the conjugacy relation is an equivalence relation Lothaire (1997). However, with the help of the following examples, we show that the $\theta$-conjugacy relation is not an equivalence relation.
@@
By definition of conjugates, if $w\in\Sigma^{+}$ and $a\in\textit{Alph}(w)$, then there exists an element in $C(w)$ that begins with $a$. However, this property does not hold for $\theta$-conjugates, as illustrated in Example 3.3. In the following, we discuss the structure of elements of the set of all $\theta$-conjugates of a word regarding this property.
@@
  1. 1.
    If $w=bccb$, then $C_{\theta}(w)=\{bccb,abcc,acbc,accb,acca\}$. Note that bccbbccb is not a $\theta$-palindrome. Now, $C_{\theta}(bcc)=\{bcc,cbc,ccb,cca\}$ and $C_{\theta}(bccb)=\{bccb\}\cup\theta(b)C_{\theta}(bcc)$.
  2. 2.
@@
For $w\in\Sigma^{+}$, $C(w)$ can contain at most $|w|$ elements and if $|C(w)|=|w|$, then each element of $C(w)$ is primitive. Mahalingam et al. (2022) proved that the set of all $\theta$-conjugates of a word $w$ can have at most $|w|+1$ elements. We now prove that for $w\in\Sigma^{*}$, if $|C_{\theta}(w)|=|w|+1$, then $C_{\theta}(w)$ must contain at least one primitive word. To prove this we need the following result.
@@
  • For $n\geq 2$, let us assume that each element of $C_{\theta}(w)$ is non-primitive. Now, the length of the root of each element in $C_{\theta}(w)$ is at most $\lfloor{\frac{|w|}{2}}\rfloor$. Consider $l=\lfloor{\frac{|w|}{2}}\rfloor$ and $w=w_{1}w_{2}\cdots w_{l}w_{l+1}\cdots w_{n}$ where each $w_{j}\in\Sigma$ for $1\leq j\leq n$. For each $i~(0\leq i\leq l-1)$, let $\alpha_{i}=\beta_{i}\gamma_{i}$ where $\beta_{i}=\theta(w_{l-i}\cdots w_{n})$ and $\gamma_{i}=w_{1}w_{2}\cdots w_{l-i-1}$. Then for each $i~(0\leq i\leq l-1)$, $|\beta_{i}|>l$ and $\alpha_{i}\in C_{\theta}(w)$. Consider $\rho(\alpha_{i})=z_{i}$ for each $i~(0\leq i\leq l-1)$. Then, $|z_{i}|\leq l$ for each $i~(0\leq i\leq l-1)$. Now, each $\alpha_{0},\alpha_{1},\cdots,\alpha_{l-1}$ has a common prefix $\beta_{0}$ where $\beta_{0}=\theta(w_{n})\theta(w_{n-1})\cdots\theta(w_{l})$ and $|\beta_{0}|>l$. This implies each $z_{i}$ is a proper prefix of $\beta_{0}$ as $|z_{i}|\leq l$ and $|\beta_{0}|>l$ where $0\leq i\leq l-1$. Since each element of $C_{\theta}(w)$ is non-primitive and $|C_{\theta}(w)|=|w|+1$, each $z_{i}$  $(0\leq i\leq l-1)$ is a distinct prefix of $\beta_{0}$. Then, the possible lengths of $z_{i}$’s are $1,2,\cdots,l$.
@@
    Otherwise, consider the length of each $z_{i}$ is distinct. Since the length of the root of each element in $C_{\theta}(w)$ is at most $l$ and there are $l$ distinct elements $\alpha_{0},\alpha_{1},\ldots,\alpha_{l-1}$ in $C_{\theta}(w)$ with a commom prefix $\beta_{0}$, we have $|\rho(\alpha_{i_{3}})|=1$ for some $0\leq i_{3}\leq l-1$. Then by Lemma 3.8, $w=a^{n}$ with $\theta(a)=a$, which is a contradiction to our assumption that $|C_{\theta}(w)|=|w|+1$.
@@
For $u,v\in\Sigma^{*}$, the equality of $C_{\theta}(u)$ and $C_{\theta}(v)$ was studied in Mahalingam et al. (2022), and the following result was proved.
@@
Mahalingam et al. (2022) Let $u,v\in\Sigma^{*}$ such that $C_{\theta}(u)=C_{\theta}(v)$.
  1. 1.
@@
    2. (b)
      $u=a^{n_{1}}u^{\prime}a^{n_{2}}$ and $v=u^{R}$ where $a\in\Sigma$, $u^{\prime}\in\Sigma^{+}$, $a\notin Pref(u^{\prime})$, $a\notin Suf(u^{\prime})$, $u^{\prime}=u^{\prime R}$, $n_{1},n_{2}\geq 1$, $n_{1}\neq n_{2}$, and $\theta(c)=c~\forall c\in\textit{Alph}(u)$.